Calculate the fluid need (in mL) of an adult weighing 88 kg who has a fever of 102.6° F. The patient is able to tolerate food and fluids
Answer #1

the fluid need per kg weight is 30ml/kg/day

so the patient's fluid need without fever = 30 × 88 =2640 ml

12% of the fluid must be increased for every 1 degree celsius rise in temperature

patient temperature is 102.6 0F or 39.2 0C

here 2 degree Celsius rise in temperature , so the fluid need of the patient per day = 2640 + 24% of the total fluid need

   = 2640+ 633.6 = 3273.6 ml

the patient needs 3274 ml per day